Consider H0: μ=72 versus H1: μ>72. A random sample of 16 observations taken from this population produced a sample mean of 75.4. The population is normally distributed with σ=6.

a. Calculate the p-value. Round your answer to four decimal places.
Answer #1

Solution :

\mu= 72

\bar x =75.4

\sigma =6

n = 16

This is the RIGHT TAIL TEST

The null and alternative hypothesis is ,

H0 :  \mu = 75

Ha : \mu > 75

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (75.4-72) /6 / \sqrt 16

= 2.27

P(z > 2.27) = 1 - P(z < 2.27 ) = 1-0.9884=0.0116

P-value = 0.0116